OpenStudy (anonymous):

Which of the following are zeros of f(x) = (x + 2) squared - 1? *check all that apply* a. x = -1 b. x = 1 c = -3 d. 3 e. none of the above

OpenStudy (anonymous):

f(x)=(x+2)^2-1 f(x)=(x+2)(x+2)-1 f(x)=x^2+4x+4-1 f(x)=x^2+4x+3 -4 +/- radical 4^2-4(1)(3) /2(1) -4 +/1 radical 16-12 /2 -4 +/ radical 4 /2 (-4+2)/2 = -1 (-4-2)/2= -3

OpenStudy (anonymous):

I think I get it... but it's both -1 & -3? you just do both + and -?

OpenStudy (anonymous):

It would be both both -3 and -1. The + and - were just there because I used the quadratic equation to solve instead of factoring.

OpenStudy (anonymous):

Ahh I see. Thank you very much!

OpenStudy (anonymous):

No problem :)